How do you find an equation in slope intercept form of the line satisfying the given condition . . . Find the coordinates of the midpoint of the segment whose endpoints are (-2, 14) and (3, -9)
Help me understand these kinds of problems please?

Answer by stanbon(75887)

How do you find an equation in slope intercept form of the line satisfying the given condition ?
Ans: That depends on the condition you are given.
If you are given two points you need to find the slope and the y-intercept.
slope = [change in y]/[change in x]
If you post a particular problem or two you will probably receive help with
this process.
-------------------
. . . Find the coordinates of the midpoint of the segment whose endpoints are (-2, 14) and (3, -9)
The x and y coordinates of the midpoint are the average of the given x and y values.
x = (2+3)/2 = 5/2
y = (14+-9)/2 = 5/2
So, the mispoint is (5/2 , 5/2)
==============
Cheers,
Stan H.